### v1

# Personalize Your Gaussian: Consistent 3D Scene Personalization from a Single Image

Abstract
Personalizing 3D scenes from a single reference image enables intuitive user-guided editing, which requires achieving both multi-view consistency across perspectives and referential consistency with the input image. However, these goals are particularly challenging due to the viewpoint bias caused by the limited perspective provided in a single image. Lacking the mechanisms to effectively expand reference information beyond the original view, existing methods of image-conditioned 3DGS personalization often suffer from this viewpoint bias and struggle to produce consistent results. Therefore, in this paper, we present Consistent Personalization for 3D Gaussian Splatting (CP-GS), a framework that progressively propagates the single-view reference appearance to novel perspectives. In particular, CP-GS integrates pre-trained image-to-3D generation and iterative LoRA fine-tuning to extract and extend the reference appearance, and finally produces faithful multi-view guidance images and the personalized 3DGS outputs through a view-consistent generation process guided by geometric cues. Extensive experiments on real-world scenes show that our CP-GS effectively mitigates the viewpoint bias, achieving high-quality personalization that significantly outperforms existing methods. The code will be released at https://github.com/Yuxuan-W/CP-GS.
Figure 1: Given a source 3DGS scene and a single reference image, CP-GS enables high-quality personalization by editing a user-specified region (e.g., the bear, man’s eye, man’s face, bench-top, entire scene) to match the reference appearance, supporting replacement, adding, and style transfer.
1 Introduction
In the evolving field of 3D computer vision, user-friendly 3D editing has attracted growing attention as a key research focus [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12]. Among recent advances, 3D Gaussian Splatting (3DGS) [13] has emerged as a groundbreaking 3D representation, offering an explicit and efficient structure that supports local manipulation and rendering in real time. Building up the 3DGS representation, we focus on a practical and intuitive form of user interaction—personalizing a 3DGS scene using only a single-view reference image—by editing a user-specified region to match the reference appearance. To make it clear, as illustrated in Figure 1, given a reference image depicting a unique brown panda, our goal is to modify a user-specific bear region in the scene to a panda that aligns to the reference appearance. This task enables intuitive 3D customization from a single image, supporting applications such as personalized avatars in virtual reality and assets stylization in interactive environments.
With the advent of large-scale pre-trained 2D diffusion models [14, 15, 16], recent 3DGS editing methods [6, 7, 12, 17] have predominately leveraged image generation models to produce pseudo-images as editing guidance that supervise the fine-tuning of 3DGS scenes. In this paradigm, the task of image-conditioned personalization requires two key consistencies in the guidance images: (1) referential consistency with the visual appearance of the reference image and (2) multi-view consistency across different perspectives to prevent conflicting guidance. However, achieving these consistencies remains a significant challenge for existing approaches [18] conditioned on a single reference image. As illustrated in Figure 2, prior methods typically adapt their image generation models directly to the single reference view, often misprojecting appearance features entangled with its geometry onto unrelated viewpoints. This leads to distorted appearances and severe multi-view inconsistencies in the editing guidance, ultimately resulting in noticeable artifacts in the final 3D output.
We argue that the core challenge lies in the viewpoint bias introduced by the limited perspective of a single reference image, where the image model lacks sufficient information to infer appearances under novel viewpoints that are far from the reference. As a result, the model is often biased towards the reference view, making existing methods struggle to produce consistent multi-view editing guidance. Therefore, in this paper, we propose Consistent Personalization for 3DGS (CP-GS), a high-quality personalization framework that addresses the viewpoint bias by progressively propagating the reference appearance to novel perspectives. As illustrated in Figure 2, to use a rough appearance as structural priors and establish viewpoint cues for guidance image generation, CP-GS operates in a coarse-to-fine manner with three stages: (1) Coarse guidance generation to initialize geometry and propagate rough appearance. (2) Iterative LoRA fine-tuning to extract and extend fine-grained reference details. (3) View-consistent generation that leverages the coarse guidance and trained LoRA to produce the final guidance images, which are used to fine-tune and produce the 3DGS output.
In the first stage, we establish a coarse guidance that serves as a structural prior, enabling the initial propagation of reference appearance into a coarse, view-consistent 3D representation. Specifically, we employ a pre-trained image-to-3D generation model [19] to produce a geometry-consistent contour with a rough texture estimate, which is integrated into the target location in the scene. As shown in Figure 2, although the resulting textures are often unrealistic due to the domain gap between the real-world reference image and the CGI-style pre-training data [20, 21] of image-to-3D model, the coarse guidance reliably captures structural geometry and a rough yet view-consistent appearance.
To recover fine-grained reference appearance free from the viewpoint bias, the second stage draws inspiration from [22], which shows that diffusion models adapted to a single image hold the potential to generate novel neighboring views around the reference. Building on this insight, we propose an iterative LoRA fine-tuning strategy that gradually extracts and propagates reference appearance to novel viewpoints. In each iteration, we translate novel-view renderings of the coarse guidance using the current model and select one well-aligned result—identified via our designed scoring mechanism based on dense feature matching [23]—to augment the training set for the next round fine-tuning.
Leveraging the coarse guidance and the trained LoRA, we employ a pre-trained flow-based model [24] in the last stage to generate the final guidance images. We begin by applying rapid rectified-flow inversion [25] to convert renderings of the coarse guidance into noisy latents, which are passed to the Flow Transformer and serve as the starting point for generation, conditioned on the depth maps of the coarse guidance. To further reduce the multi-view inconsistency arising from viewpoint variance, we introduce an epipolar-constrained token replacement strategy that aggregates visual features across all views based on geometric correspondences, effectively improving overall multi-view coherence.
As illustrated in Figure 1, by progressively propagating the single-view reference appearance in a coarse-to-fine manner, CP-GS effectively addresses the challenges posed by the viewpoint bias, resulting in superior visual quality in personalized 3DGS results. Comprehensive evaluations across diverse real-world scenes demonstrate that our CP-GS successfully address the artifacts caused by limited reference perspectives and outperforms state-of-the-art methods in both qualitative and quantitative comparisons. Based on the above, our contributions can be summarized in three aspects:
  • We identify the viewpoint bias caused by limited reference perspective information as the crux of referential and multi-view inconsistencies in previous single-view 3D personalization methods.
  • To mitigating the viewpoint bias, we propose a coarse-to-fine appearance propagation framework that progressively expands the single-view reference appearance to novel perspectives, generating guidance images with faithful referential consistency and strong multi-view consistency.
  • We validate CP-GS through extensive experiments on various real-world scenes, demonstrating its superior performance over previous 3DGS personalization and editing methods in both qualitative and quantitative evaluations.
2 Related Works
Figure 2: red: Previous methods suffer from viewpoint bias and produce distorted editing guidance, leading to both the referential and multi-view inconsistencies. blue: By progressively propagating reference to novel views, CP-GS mitigates the bias and achieves both consistencies in the guidance.
Image-guided 2D Customization. Given a set of reference images, the task of 2D customization aims to edit a source image or generate a new image under the guidance of the reference, where customization methods [26, 27, 28] typically optimize a special token or use LoRA-based adaptation to capture the appearance of the reference images. Built on this strategy, early methods [26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36] rely on multiple reference images to construct the novel content through test-time fine-tuning (TTF). Subsequent works [37, 38, 39, 40] further improve the flexibility of this paradigm by training with a single reference image. Recently, leveraging large-scale image datasets [41, 42, 43, 44, 45, 46], a line of work [47, 48, 49, 50, 51, 52, 53] has adopted pre-trained adaptation (PTA), which trains on large-scale paired data and bypasses fine-tuning during inference. While our iterative LoRA fine-tuning strategy builds on the test-time fine-tuning paradigm, the task of 3D personalization presents additional challenges beyond those in 2D customization, notably the need for multi-view consistency and mitigating the viewpoint bias.
Consistent 3D Field Editing. Early approaches for consistent 3D editing [54, 55, 56, 57, 58, 59, 60, 61, 62, 63] predominantly rely on NeRF [64] representations optimized via Score Distillation Sampling (SDS)-based techniques [65]. Subsequent works [1, 2, 4] employ image-guided 3D editing by leveraging pre-trained 2D diffusion models to generate multi-view guidance images. Pioneered by [1, 12], recent methods integrate Gaussian Splatting [13] into 3D field editing due to its superior efficiency and controllability. More recently, a line of research [6, 7, 8, 17] has aimed to explicitly ensure multi-view consistency in the guidance images. VcEdit [7] introduces latent and attention map aggregation, while GaussCtrl [17] and DGE [6] utilize cross-view extensive attention to harmonize the variations across views. However, all these methods are limited to simple text prompts condition and lack the ability of customized editing. The most relevant work with ours is TIP-Editor [18], which combines LoRA and SDS to distill the reference content into 3D scene. However, it fails to consistently expand the reference appearance across views, often exhibits visual artifacts in the 3D outputs due to viewpoint bias.
3 Methodology
Figure 3: The pipeline of our CP-GS includes three stages: coarse guidance generation via a pre-trained image-to-3D model; iterative LoRA fine-tuning to extract and propagate detailed reference appearance; and view-consistent generation of guidance images to produce final 3DGS outputs.
In this section, we present the CP-GS personalization framework from a single-view reference image (Sec. 3.1), with the overall pipeline illustrated in Figure 3. We first employ a pre-trained image-to-3D model to construct a coarse guidance with rough yet view-consistent reference appearance, serving as the initial step of our propagation (Sec. 3.2). To further extract and propagate fine-grained reference appearance, we then introduce an iterative LoRA fine-tuning strategy that progressively expands the training views through image translation and selective augmentation (Sec. 3.3). Finally, we combine the coarse guidance and the trained LoRA within a pre-trained Flow model to generate multi-view consistent, reference-aligned guidance images, resulting in the final 3DGS output (Sec. 3.4).
3.1 Problem Definition
Given a source 3DGS scene $\mathcal{G}^{\text{src}}$ and a reference image $\mathcal{I}^{\text{ref}}$, the goal is to edit a user-specific region in $\mathcal{G}^{\text{src}}$ to the personalized $\mathcal{G}^{\text{edit}}$ that align with $\mathcal{I}^{\text{ref}}$. To achieve this, we adopt an image-guided paradigm that generates a set of multi-view personalized guidance images $\mathcal{I}^{\text{gde}}$ to supervise the transformation of $\mathcal{G}^{\text{src}}$ into the output $\mathcal{G}^{\text{edit}}$. We define an editing loss for each view by combining a mean absolute error $\mathcal{L}_{\text{MAE}}$ and a perceptual loss $\mathcal{L}_{\text{LPIPS}}$ between the real-time rendering and corresponding guidance image. The final 3DGS model $\mathcal{G}^{\text{edit}}$ is optimized by minimizing the total loss across all views $\mathcal{V}$:
$\mathcal{G}^{\text{edit}}=\>\underset{\mathcal{G}}{\mathrm{argmin}}\>\sum_{v\in\mathcal{V}}{(\lambda_{1}\mathcal{\mathcal{L}_{\text{MAE}}}(\mathcal{R}(\mathcal{G},v),\>\mathcal{I}^{\text{gde}})+\lambda_{2}\mathcal{\mathcal{L}_{\text{LPIPS}}}(\mathcal{R}(\mathcal{G},v),\>\mathcal{I}^{\text{gde}}))},$ (1)
where $\mathcal{R}$ denotes the rendering function [13]. This paradigm requires the multi-view guidance images $\mathcal{I}^{\text{gde}}$ to satisfy two key properties: multi-view consistency across $\mathcal{V}$ to prevent optimization conflicts, and referential appearance consistency to the $\mathcal{I}^{\text{ref}}$ to fulfill the personalization objective. Our CP-GS is designed to explicitly ensures both consistency to achieve high-quality 3D personalization.
3.2 Coarse Guidance Generation
As noted in Sec. 1, the limited perspective of a single reference image fails to provide sufficient geometric and coherent appearance information for constructing a consistent multi-view representation. Therefore, in the first stage, we leverage an off-the-shelf image-to-3D generation model TRELLIS [19], pre-trained on large-scale CGI-style 3D datasets [20, 21], to produce a coarse guidance scene that expands the reference into a rough yet multi-view consistent representation. As illustrated in the top-left of Figure 3, the reference image is fed into the pre-trained TRELLIS to generate the corresponding 3D rough asset, which is then integrated into the source scene to replace or augment the user-specific target region. Here we provide two integration modes: (1) Adding new content – the user provides a 3D bounding box specifying the object’s position and scale; (2) Replacing existing content – the target bounding box is extract from the existing content via PCA [66], and the generated asset is fitted accordingly. This coarse guidance provides a plausible 3D geometry and establishing a rough yet view-consistent appearance that serves as a structural prior for subsequent stages.
Figure 4: (a) Visualization of the translated results and the corresponding similarities under our scoring mechanism. (b) Illustration of the proposed epipolar-constrained token replacement strategy.
3.3 Iterative LoRA Fine-tuning
Due to the inevitable domain gap between the real-world reference image and the CGI-style datasets [20, 21] used to train the image-to-3D model, the generated assets in coarse guidance often exhibit unrealistic and rough appearance that lacks referential consistency. Under the single image setting, we observe that image generation model also tends to overfit to the reference perspective, resulting in a strong viewpoint bias. To address these issues, we adopt an iterative LoRA [67] fine-tuning strategy that retrieves a fine-grained appearance from the reference and progressively propagates it to novel views.
Specifically, we initialize the LoRA training set with the given single-view image, conducting the first iteration of fine-tuning using a prompt containing a special token to encode the reference characteristics. Inspired by DreamBooth3D [22], which demonstrates that image generation models [14] adapted to a single image can synthesize novel views of the reference subject within a limited range around the training perspective, we render the coarse guidance from multiple viewpoints and apply the fine-tuned model to translate their appearance toward the reference target using the same prompt. Subsequently, we select one well-aligned translated image using a task-specific scoring mechanism and append it to the training set to augment the next round of fine-tuning.
We notice that designing such scoring mechanism is non-trivial, as it must avoid viewpoint-biased translations and redundant views that are already well covered by the training set, ensuring that each selection contributes meaningfully to appearance propagation. Notably, as shown in Figure 4(a), both types of undesirable cases tend to exhibit high similarity to the training images: (1) redundant views, which are close to the training perspective, naturally share similar appearance; and (2) biased translations, which often inherit excessive training-view features due to overfitting, also tend to exhibit higher similarity to training images than the well-aligned novel-view results. Therefore, we identify the well-aligned result as the one with minimal overall similarity to the training set, measured via dense feature matching using the pre-trained RoMa model [23]. Denoting $\mathcal{I}^{\text{train}}_{t}$ the training image set and $\mathcal{I}^{\text{trans}}_{t}$ the translations at iteration $t$, our scoring and selection are formulated as:
$\mathcal{I}^{\text{train}}_{t+1}=\mathcal{I}^{\text{train}}_{t}\cup\{\operatorname*{arg\,min}_{i}\sum_{j}\mathbf{S}_{\text{RoMa}}(I^{\text{trans}}_{i},I^{\text{train}}_{j})\},\>\>\text{where}\>\>I^{\text{trans}}_{i}\in\mathcal{I}^{\text{trans}}_{t},\>\>I^{\text{train}}_{j}\in\mathcal{I}^{\text{train}}_{t}$ (2)
where $\mathbf{S}_{\text{RoMa}}(\cdot)$ denotes the similarity computed by RoMa [23] model. Leveraging the neighboring-view generation capability of the LoRA module, our iterative fine-tuning strategy effectively propagates single-view reference details to novel perspectives and alleviates the viewpoint bias, enabling the model to learn fine-grained appearance with both multi-view and reference consistency.
Figure 5: Additional personalization result of our CP-GS, demonstrating high-quality 3DGS scene customization that faithfully align with the reference image across various scenarios.
Figure 6: Qualitative comparison of personalization results between our CP-GS and the existing SOTA methods [6, 18], where CP-GS outperforms with superior visual quality and reference alignment.
3.4 View-consistent Generation
In the final stage, we adopt a view-consistent generation straregy based on a pre-trained Flow Transformer [24], combining the coarse guidance (Sec. 3.2) and the iteratively trained LoRA module (Sec. 3.3) to produce the final consistent guidance images. As shown in Figure 3, we begin by rendering the coarse guidance into multi-view images and converting them into noisy latents using rectified-flow inversion [25] to encode both the appearance and geometry. Serving as the starting point for subsequent generation process, these latents are then fed into the Flow Transformer along with rendered depth maps, which provide geometric cues to align appearance generation with the underlying structure and enhance multi-view consistency.
Inspired by [6, 68], we introduce an epipolar-constrained token replacement mechanism, to promote multi-view consistency by unifying foreground tokens across views that correspond to the same 3D locations. We perform token replacement in the early dual-stream blocks of the Flow Transformer, where visual tokens are explicitly maintained and can be directly modified. During generation, we automatically select a set of key views with minimal overlap to ensure full coverage, and extract foreground pixel indices in all viewpoints using masks from the coarse guidance. As illustrated in Figure 4(b), for each foreground token in non-key views, we compute its epipolar lines on the two nearest key views and replace it with an interpolation of the most similar tokens along those epipolar lines, weighted by the camera distance to each key view. Given a non-key frame, the interpolated token $\mathbf{f}^{\prime}(\mathbf{u})$ at pixel $\mathbf{u}$ used to replace the original token $\mathbf{f}(\mathbf{u})$ is computed using the foreground tokens of the two nearest key frames $\mathbf{k}_{i}$, indexed by $i\in\{1,2\}$. Letting $c$ denote the non-key camera and $l_{\mathbf{u}\rightarrow i}$ denote the epipolar line of $\mathbf{u}$ in each key view, the token $\mathbf{f}\textquoteright(\mathbf{u})$ is computed as:
$\mathbf{f}^{\prime}(\mathbf{u})=\sum_{i}\mathbf{k}_{i}(\mathbf{v}_{i})\mathcal{D}(c,c_{i})/\sum_{i}\mathcal{D}(c,c_{i}),\>\>\text{where}\>\>\mathbf{v}_{i}=\underset{\mathbf{v}\in l_{\mathbf{u}\rightarrow i}}{\operatorname*{arg\,max}}\>\langle\mathbf{f}(\mathbf{u}),\mathbf{k}_{i}(\mathbf{v})\rangle$ (3)
where $\mathcal{D}(c,c_{i})$ represents the camera distance from $c$ to each key view’s camera $c_{i}$. This mechanism effectively alleviates cross-view variance, producing fine-grained guidance images with multi-view consistency and faithful reference alignment. These images then supervise the 3DGS parameter updating of the coarse guidance, yielding the final personalized 3DGS result of our CP-GS framework.
4 Experiments
4.1 Implementation Details
We implement our framework based on the official 3DGS codebase [13], GaussianEditor [1], and the LoRA training scripts from Diffusers [69]. We employs TRELLIS [19] as image-to-3D model to generate our coarse guidance. For the Flow transformer, we adopt FLUX.1-dev [24] equipped with the depth LoRA adapter. In most cases, we use two iterations of LoRA training to propagate the appearance, which takes around 15 minutes on two NVIDIA RTX A6000 GPUs and is reusable across different source scenes. Using the trained LoRA and coarse guidance, we generate consistent multi-view guidance images and optimize the 3DGS model with the Adam optimizer [70] at a learning rate of 0.001, taking around 10 minutes per scene when using the same two A6000 GPUs.
4.2 Qualitative Evaluation
We compare our CP-GS with two state-of-the-art 3DGS editing baselines: DGE [6], which conditioned on text prompt, and TIP-Editor [18], the only existing method with the same single-image condition as ours. We construct a challenging test set comprising reference images from TIP-Editor and additional internet-sourced examples with highly specialized, visually intricate appearances. For DGE, we employ GPT-4o [71] to generate concise captions (within 5 words) that describe the reference object, as longer prompts were observed to degrade its performance. As shown in Figure 6, both baselines fail to preserve the distinctive appearance features of the reference images. Moreover, TIP-Editor exhibits severe artifacts in its personalized results, primarily due to multi-view inconsistencies in the guidance images resulting from viewpoint bias. In contrast, our CP-GS consistently produces clean, coherent, and intricately detailed edits that faithfully align with the reference image.
This performance gap underscores the inability of the baseline methods to capture and propagate reference appearance. In particular, DGE illustrates the shortcomings of text-conditioned 3DGS editing for personalization: lacking direct access to the reference image, it relies solely on short textual prompts that fail to capture rich visual details. Moreover, the specialized reference images often fall outside the distribution of text-to-image models, making them difficult to represent accurately. On the other hands, TIP-Editor lacks explicit mechanism to extend the reference appearance to novel viewpoints, resulting in strong viewpoint bias, which introduces multi-view inconsistencies in its 2D guidance, ultimately leading to visual artifacts in the 3DGS results. By contrast, CP-GS explicitly addresses these issues through the coarse-to-fine appearance propagation, enabling high-quality 3DGS personalization that ensures both referential and multi-view consistency. Additional results showcasing the effectiveness of our CP-GS are presented in Figure 5 and the Appendix.
Figure 7: Comparison between the coarse guidance and our final results, where our final results effectively refine the unrealistic and visually discordant appearance presented in the coarse guidance.
Table 1: Quantitative comparison between our CP-GS and the existing SOTA methods [6, 18], where CP-GS significantly outperform others in both visual quality and the alignment with reference image.
[…]
DGE [6] 31.89% 6.37% 41.73 67.26 14.22
TIP-Editor [18] 25.46% 17.28% 43.88 70.92 14.46
CP-GS (Ours) 78.28% 80.09% 50.33 76.78 18.03
Figure 8: Ablation comparison on the guidance images produced by three specific configurations: training the LoRA module only on the single-view reference image (Sing. LoRA), using our iterative LoRA fine-tuning yet excluding the constrained token replacement (Iter. LoRA), and our Full Version.
4.3 Quantitative Evaluation
In Table 1, we present a quantitative evaluation comparing our CP-GS with the two baselines [6, 18] on over 20 samples collected in the same manner as the qualitative experiments. We first conduct a user study to assess the proportion of results deemed satisfactory by users in aspects of visual quality and the reference alignment, with further details provided in the Appendix. Besides, we follow existing setting [18] to report CLIP [73] and DINO [72] image-to-image similarity metrics that quantify the alignment between edited outputs and the reference image by computing visual feature similarity. In addition, we adopt the CLIP directional similarity [74] to measure the semantic alignment between the text prompt and the semantic shift from the source to edited results. As shown in Table 1, our CP-GS consistently outperforms all baselines across both perceptual metrics and user study evaluations. This performance gap stems from two main limitations in these baselines: (1) DGE lacks direct access to the reference image and relies solely on short text prompts, which fail to capture fine-grained appearance details; (2) TIP-Editor fails to propagate the reference appearance to novel views, resulting in strong viewpoint bias that introduces multi-view conflicts in the editing guidance and ultimately leads to visual artifacts and poor reference alignment. These results highlight the superior performance of our CP-GS, underscoring the critical role of capturing and expanding reference appearance across novel viewpoints in tackling single-view conditioned 3DGS personalization.
4.4 Ablation Study
To analysis the contribution of each component, in this section, we compare the coarse guidance with our final results and conduct ablation studies on the iterative LoRA fine-tuning strategy and epipolar-constrained token replacement. Additional quantitative ablation study is provided in Appendix.
Coarse Guidance vs. Final Results. In the first stage of our method, a coarse asset is generated by the image-to-3D model (Sec. 3.2) and integrated into the source scene to produce the coarse guidance. A natural question arises: can this coarse guidance suffice as the final edit? To assess the necessity of our subsequent stages, we compare the coarse guidance with our final personalization results in Figure 7. This comparison reveals two major limitations of the coarse guidance: (1) The domain gap between real-world reference images and the CGI-style training data of the 3D generation model [19] results in overly smooth and grid-like unrealistic textures that fail to reflect the photorealic appearance of the reference. (2) Direct insertion of the generated asset leads to poor contextual blending, where the inserted object often appears visually detached from the 3DGS scene (e.g., the sunglasses example), especially around boundaries. In contrast, our final results exhibit rich, realistic textures that closely resemble the reference image and blend seamlessly with the source 3DGS scene, demonstrating the effectiveness and necessity of the subsequent refinement stages.
Iterative LoRA Fine-tuning. We compare two configurations to assess the contribution of our iterative LoRA fine-tuning: using only the single-view reference image for fine-tuning (Sing. LoRA) versus applying our iterative expansion strategy (Iter. LoRA) introduced in Sec. 3.3. The top 2 rows in Figure 8 shows the resulting guidance images from these variants. In the presented viewpoints that deviate from the reference, Sing. LoRA misprojects the appearance entangled with reference-view geometry to unrelated views, resulting in noticeable distortion and multi-view inconsistency. In contrast, Iter. LoRA significantly alleviates this distortion, generating appearance correctly adapted to these novel perspectives. This highlights the importance of progressively expanding the reference coverage in mitigating the viewpoint bias and producing multi-view consistent guidance images.
Constrained Token Replacement. To further reduce the cross-view variance during generation, we adopt an epipolar-constrained token replacement strategy during the generation of final guidance images (Sec. 3.4). The bottom 2 rows in Figure 8 compares the Full Version that includes this mechanism and the Iter. LoRA variant where it is disabled. Although Iter. LoRA successfully expands the reference appearance to novel viewpoints and resolves major distortions, it still suffers from subtle multi-view inconsistencies in visual details (e.g., the mouth orientation of the dog, and the doll’s eyelashes). In contrast, our Full Version leverages 3D-aware token replacement guided by epipolar constraints and eliminate such inconsistencies, producing guidance images with improved multi-view appearance consistency that enhance the coherence and visual fidelity of the final 3DGS results.
5 Conclusion
In this paper, we presented CP-GS, a novel framework for consistent and personalized 3D scene editing from a single-view reference image. To address the visual artifacts in existing image-conditioned methods caused by viewpoint bias and limited reference perspective, CP-GS introduces a coarse-to-fine reference propagation framework that integrates coarse guidance generation, iterative LoRA fine-tuning, and a view-consistent generation stage leveraging geometric cues and epipolar-constrained token replacement. These components enable the generation of guidance images with strong multi-view consistency and faithful referential consistency, producing high-quality 3DGS results. Extensive experiments across real-world scenes demonstrate that CP-GS markedly outperforms existing methods in both visual quality and reference alignment, enabling high-quality 3DGS personalization for real-world applications. In the future, we aim to improve CP-GS’s efficiency by distilling it into a single-pass pipeline and enhancing robustness to occluded reference images.